Please help me solve this problem: The greater of two consecutive even integers is twelve less than twice the smaller.Find the consecutive even integers.
Answer by nerdybill(7384)

Please help me solve this problem: The greater of two consecutive even integers is twelve less than twice the smaller.Find the consecutive even integers.
.
Let x = smaller of two consecutive even integers
then
x+2 = larger of two consecutive even integers
.
From: "The greater of two consecutive even integers is twelve less than twice the smaller." we get:
x+2 = 2x-12
2 = x-12
14 = x (smaller integer)
.
larger integer:
x+2 = 14+2 = 16 (larger integer)
.
Solution: 14 and 16